# Spectral gap of a convex combination of a random permutation and a deterministic matrix

### v3

@@ 4.3 Freeness with amalgamation over $\mathcal{B}$ of $u$ and $\mathcal{M}$
    $u^{\alpha_{m}}b_{m}u^{\alpha_{m-1}}\cdots u^{\alpha_{1}}b_{1}u^{\alpha_{0}}e(k)=\#e(v)\text{ and }\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ |v|=\sum_{i=1}^{m}|\alpha_{i}|$ (17)
    with #$\#\in\{0,1\}$. The vertex $v$ verifies $v=(v^{-},(\epsilon(\alpha_{m}),j_{\epsilon(\alpha_{m})}(v^{-})))$ for some $v^{-}\in\mathcal{V}$.
@@
[…]
@@
[…]
@@
We can now prove 4.3.
Proof of 4.3.
For the first point we proceed by induction over $|v|\geq 1$. Let $v:=(k,i_{1})=(k,(\epsilon,k_{1}))$. Either $k_{1}=j_{\epsilon}(k)$, then $e(v)=e(k)\otimes e(\epsilon,j_{\epsilon}(k))=u^{\epsilon}(e(k))$, or $j_{\epsilon}(k)\neq k_{1}$ and then denoting $B:=E_{j_{\epsilon}(k),k_{1}}$ and $b=\Pi(B)$ we have that $bu^{\epsilon}(e)=b(e(k)\otimes e(j_{\epsilon}(k),\epsilon))=e(k)\otimes Be(\epsilon,j_{\epsilon}(k))=e(v)$. We now consider $v=(k,i_{1},...,i_{p+1})$ for some $p\geq 1$. For $t<p+1$ we consider $v_{t}:=(k,i_{1},...,i_{t})$ and $e_{t}:=e(v_{t})$. Since $e=e(v)\in\Xi$, we have $v_{t}\in\mathcal{V}$ for all $t<p+1$. Let $\alpha_{0},...,\alpha_{m}\in\mathbb{Z}$ and $b_{1},...,b_{m}$ as described in 4.3 and such that
@@
We now prove the second assertion and we start with the case where we impose also $\alpha_{m}\neq 0$. We fix $k\in[N]$ and proceed by induction, this time on $p:=\sum_{i=0}^{m}|\alpha_{i}|$. For $p=1$ we necessarily have that $m=1$ and $\alpha_{1}=\alpha\in\{+1,-1\}$. For $B=E_{ij}$ an elementary matrix, we have either $j\neq k$ and then $u^{\alpha}be(k)=be(k)=0$, or $j=k$ and then $u^{\alpha}be(k)=u^{\alpha}e(i)=e(i)\otimes e(\alpha,j_{\alpha}(i))$. We now consider $p\geq 2$. If $m=1$ and $|\alpha_{1}|=m$ then we conclude with 4.4. Otherwise we set
@@
and $\epsilon_{i}:=\operatorname{sign}(\alpha_{i})$. Since $\alpha_{m}\neq 0$ then by induction we have $v\in\mathcal{V}$ such that we have $f=e(v)=e(v^{-})\otimes e(\epsilon_{m-1},j_{\epsilon_{m-1}}(v^{-}))$. We consider $E_{ij}$ the elementary matrix such that $b_{m}=\Pi(E_{ij})$. If $j_{\epsilon_{m-1}}(v^{-})\neq j$ then $b_{m}f=0$ and therefore $e=0$. Otherwise we have that $b_{m}f=e(v^{-})\otimes e(\epsilon_{m-1},i)$ with $i\neq j_{\epsilon_{m-1}}(v^{-})$. Applying 4.4 to $e=b_{m}f$ and $\alpha=\alpha_{m}$ we obtain the expected result. Now if $\alpha_{m}=0$, knowing that $\alpha_{m-1}\neq 0$ we find $v\in\mathcal{V}$ of length $p=\sum|\alpha_{i}|$ such that we have $u^{\alpha_{m-1}}b_{m-1}\cdots u^{\alpha_{0}}e(k)=\epsilon e(v)$. Applying $b_{m}$ to $e(v)$ we either find 00 or change the last coordinate $i_{p}$ of $v$.
We now prove the freeness with amalgamation of $u$ and $\mathcal{M}$, i.e. that (16) holds true for all $e\in\Xi$ and for all $b_{i}=\Pi(B_{i})$ where $B_{i}\in\mathrm{M}_{N}(\mathbb{C})$ and $\mathrm{E}(B_{i})=0$. Owing to the first assertion of 4.3, we only have to check (16) holds for $e\in H_{\emptyset}$ basis vector. Finally by linearity in $b_{i}$ for all $i\leq n$ it is sufficient to consider that $B_{i}$ is an elementary matrix $E_{k,\ell}$ of $\mathrm{M}_{N}(\mathbb{C})$ with zeros on the diagonal. Now let $k\in[N]$, $m\geq 2$, $\alpha_{0},...,\alpha_{m}\in\mathbb{Z}$ such that $\alpha_{i}\neq 0$ for all $i\notin\{0,m\}$ and $b_{1},...,b_{m}$ such that $b_{i}=\Pi(B_{i})$ where $B_{i}$ elementary matrix. Owing to the second assertion in 4.3, we have $f=u^{\alpha_{m}}b_{m}\cdots u^{\alpha_{1}}b_{1}u^{\alpha_{0}}e(k)=\epsilon e(v)$ for some $v\in\mathcal{V}$ such that $|v|\geq 1$. In particular $f$ is indeed orthogonal to $e$. ∎
@@ 4.4 Spectral radius of the limit operator
$\mathcal{L}(T,n):=\left\{\vec{n}=(n_{1},...,n_{T})\in\{0,...,n\}^{T}:\leavevmode\nobreak\ \sum_{t=1}^{T}n_{t}+T-1=n\right\}.$
We first express $\Phi[a^{n_{1}}(a^{*})^{n_{2}}]$ and $\Phi[p^{n_{1}}(p^{*})^{n_{2}}]$ for any integer $n_{1},n_{2}\geq 0$ and $a,p$ defined by (10).
@@
$\Phi[a^{n_{1}}(a^{*})^{n_{2}}]=\sum_{T=1}^{n_{1}\land n_{2}}\sum_{\begin{subarray}{c}\vec{n}\in\mathcal{L}(T,n_{1})\\ \leavevmode\nobreak\ \vec{n}^{\prime}\in\mathcal{L}(T,n_{2})\leavevmode\nobreak\ \end{subarray}}\prod_{t=1}^{T}\operatorname{Tr}_{N}\left(Q^{n_{t}}(Q^{*})^{n_{t}^{\prime}}\right)$
@@
$\Phi[p^{n_{1}}(p^{*})^{n_{2}}]=\sum_{T=1}^{n_{1}\land n_{2}}\sum_{\begin{subarray}{c}\vec{n}\in\mathcal{L}(T,n_{1})\\ \leavevmode\nobreak\ \vec{n}^{\prime}\in\mathcal{L}(T,n_{2})\leavevmode\nobreak\ \end{subarray}}(1-r)^{2(T-1)}\prod_{t=1}^{T}\operatorname{Tr}_{N}\left(\{rQ\}^{n_{t}}\{rQ^{*}\}^{n_{t}^{\prime}}\right).$
@@
$\mathcal{F}_{t}:=\sigma\left(\{\alpha\},\{j_{\epsilon}(s)\leavevmode\nobreak\ ;s\in\mathcal{S},\leavevmode\nobreak\ |s|\leq t-1,\leavevmode\nobreak\ \epsilon\in\{+,-\}\}\right).$
Then 4.5 is a direct consequence of the following lemma.
@@
According to the first assertion of the lemma, we find a unique $(w_{T-1},w^{\prime}_{T-1})\in\mathcal{V}^{2}$ of length $T-1$ such that $e_{T-1}=e(w_{T-1})$ and $e^{\prime}_{T-1}=e(w_{T-1}^{\prime})$. Then rewriting $I_{T}$ and supposing that $w_{T-1}$ and $w_{T-1}^{\prime}$ are non-zero and therefore $\epsilon(w_{T-1})=\epsilon(w_{T-1}^{\prime})=+$ we obtain
@@
where for $M\in\mathrm{M}_{N}(\mathbb{C})$ we use the notation $(M)_{k}:=(M)_{k,k}$. Noticing that for all $w\in\mathcal{V}$ such that $|w|=T-1$ we have that $j_{+}(w)$ is independent from $\mathcal{F}_{T-1}$ we have
$\mathbb{E}(I_{T}|\mathcal{F}_{T-1})=\sum_{w\in\mathcal{V}\leavevmode\nobreak\ ;|w|=T-1}\mathbb{E}((b_{T}c_{T})_{j_{+}(w)})\mathds{1}(w_{T-1}=w_{T-1}^{\prime}=w)=\operatorname{Tr}_{N}(b_{T}c_{T})\langle e(w_{T-1}),e(w_{T-1}^{\prime})\rangle.$
@@
For all $B\in\mathrm{M}_{N}(\mathbb{C})$ and $n\geq 1$ integer we have:
@@
Let $e\in\mathrm{H}$ such that $\|e\|_{2}=1$. We set $b:=\Pi(B)$. We have:
$[u+b]^{n}(e)=u^{n}(e)+\sum_{T=0}^{n}\sum_{\sum_{t=1}^{T}\lambda_{t}=n-T+1}b^{\lambda_{1}}ub^{\lambda_{2}}u\cdots ub^{\lambda_{T}}$
where in the term $b^{\lambda_{1}}ub^{\lambda_{2}}u\cdots ub^{\lambda_{T}}$, the number of occurrences of $u$ is bounded by $n-1$, therefore the sum is orthogonal to $u^{n}(e)$. It remains to take the norm of in the equality above. ∎
5 Main technical statement
Having proved the first statement of 2.2 about the existence of the operator $(q,u)\in\mathcal{B}(\mathrm{H})$, the second statement is proved by applying the following theorem. First note that for any $\ell\geq 0$ integer we have
$\rho(A_{\#}|_{\boldsymbol{1}^{\perp}})=\{\rho(A_{\#}^{\ell}|_{\boldsymbol{1}^{\perp}})\}^{1/\ell}\leq\|A_{\#}|_{\boldsymbol{1}^{\perp}}^{\ell}\|^{1/\ell},$
To simplify further computations, we introduce
$A_{\ell,\#}:=MA_{\#}^{\ell}$
for any $\ell\geq 1$ integer. Since $M$ is a permutation and $Q$ verifies (H:1), $A_{\ell,\#}$ and $A_{\#}^{\ell}$ both leave $\boldsymbol{1}$ and $\boldsymbol{1}^{\perp}$ stable and have the same operator norm. Furthermore for $N$ large enough we have
@@ 5 Main technical statement
Where $\epsilon=\epsilon(N)$ is given by 2.2. Therefore for $N$ large enough
$\mathbb{P}\left(\|A^{\ell}|_{\boldsymbol{1}^{\perp}}\|\geq(1+\epsilon)^{\ell}\rho_{n}^{\ell}\right)=\mathbb{P}\left(\|A_{\ell}|_{\mathbf{1}^{\perp}}\|\geq(1+\epsilon)^{\ell}\rho_{n}^{\ell}\right)\leq\mathbb{P}\left(\|{A_{\ell}|}_{\boldsymbol{1}^{\perp}}\|\geq e^{\frac{\epsilon\ell}{2}}\rho_{n}^{\ell}\right).$
The same reasoning holds replacing $A$ with $A_{r}$ and $\rho_{n}$ with $\rho_{n,r}$. Adjusting the constants, the second statement of 2.2 becomes a corollary of the following result applied to $\ell=\frac{(1-c)\ln(N)}{7K_{\#}^{\prime}\ln(d)}$ where
$K^{\prime}=1\vee(2\ln(1+\delta+\kappa))\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ K^{\prime}_{r}=1\vee\left(2\ln\left(\frac{1+r\delta+r\kappa}{1-r}\right)\right).$
The condition (H:2) implies that $K_{\#}^{\prime}$ is uniformly bounded in $N$.
@@
Let $n\geq 0$, $(M,Q)$ and $\rho_{n,\#}$ be as in 2.2. For any $10<c<$ 1, there exists $C>0$ depending on $n$, $K$ and $\alpha$, such that for any $1\leq\ell\leq\frac{\ln(N)}{7\ln(d)}$ we have:
$\mathbb{P}\left(\|A_{\#}^{\ell}|_{\boldsymbol{1}^{\perp}}\|\geq e^{C\ln(N)^{1-\alpha}}\rho_{n,\#}^{\ell}\right)\leq N^{-c}+C\frac{\ell^{4}(4d)^{3\ell}}{N}.$
To estimate an upper bound on $\|(A_{\ell,\#})|_{\boldsymbol{1}^{\perp}}\|$, we use the high trace method, i.e. Markov inequality combined with the inequality

### v4

@@ 6.3 Freeness with amalgamation over $\mathcal{B}$ of $u$ and $\mathcal{M}$
    $u^{\alpha_{m}}b_{m}u^{\alpha_{m-1}}\cdots u^{\alpha_{1}}b_{1}u^{\alpha_{0}}e(k)=\#e(v)\text{ and }\penalty 10000\ \penalty 10000\ \penalty 10000\ |v|=\sum_{i=1}^{m}|\alpha_{i}|$ (18)
    with $\in\{0,1\}$. The vertex $v$ verifies $v=(v^{-},(\epsilon(\alpha_{m}),j_{\epsilon(\alpha_{m})}(v^{-})))$ for some $v^{-}\in\mathcal{V}$.
@@
Point 2 of the previous lemma says that the length of $s\in\mathcal{S}$ is the distance between $s$ and the root $\{\emptyset\}$. Besides one can notice that the operator $b=\Pi(B)$, with $B$ an elementary matrix with zeros on the diagonal is the adjacency operator of the graph with only one edge in each copie of $[N]$.
@@
$\text{ for }1\leq t\leq|\alpha|,\penalty 10000\ v_{\epsilon(t+1)}$ $:=(v_{\epsilon t},(\epsilon,j_{\epsilon}(v_{\epsilon t})))\in\mathcal{V}.$
@@
We can now prove ˜6.1.
Proof of ˜6.1.
For the first point we proceed by induction over $|v|\geq 1$. Let $v:=(k,i_{1})=(k,(\epsilon,k_{1}))$. Either $k_{1}=j_{\epsilon}(k)$, then $e(v)=e(k)\otimes e(\epsilon,j_{\epsilon}(k))=u^{\epsilon}(e(k))$, or $j_{\epsilon}(k)\neq k_{1}$ and then denoting $B:=E_{j_{\epsilon}(k),k_{1}}$ and $b=\Pi(B)$ we have that $bu^{\epsilon}(e)=b(e(k)\otimes e(j_{\epsilon}(k),\epsilon))=e(k)\otimes Be(\epsilon,j_{\epsilon}(k))=e(v)$. We now consider $v=(k,i_{1},...,i_{p+1})$ for some $p\geq 1$. For $t<p+1$ we consider $v_{t}:=(k,i_{1},...,i_{t})$ and $e_{t}:=e(v_{t})$. Since $e=e(v)\in\Xi$, we have $v_{t}\in\mathcal{V}$ for all $t<p+1$. Let $\alpha_{0},...,\alpha_{m}\in\mathbb{Z}$ and $b_{1},...,b_{m}$ as described in ˜6.1 and such that
@@
We now prove the second assertion and we start with the case where we impose also $\alpha_{m}\neq 0$. We fix $k\in[N]$ and proceed by induction, this time on $p:=\sum_{i=0}^{m}|\alpha_{i}|$. For $p=1$ we necessarily have that $m=1$ and $\alpha_{1}=\alpha\in\{+1,-1\}$. For $B=E_{ij}$ an elementary matrix, we have either $j\neq k$ and then $u^{\alpha}be(k)=be(k)=0$, or $j=k$ and then $u^{\alpha}be(k)=u^{\alpha}e(i)=e(i)\otimes e(\alpha,j_{\alpha}(i))$. We now consider $p\geq 2$. If $m=1$ and $|\alpha_{1}|=m$ then we conclude with ˜6.2. Otherwise we set
@@
and $\epsilon_{i}:=\operatorname{sign}(\alpha_{i})$. Since $\alpha_{m}\neq 0$ then by induction we have $v\in\mathcal{V}$ such that we have $f=e(v)=e(v^{-})\otimes e(\epsilon_{m-1},j_{\epsilon_{m-1}}(v^{-}))$. We consider $E_{ij}$ the elementary matrix such that $b_{m}=\Pi(E_{ij})$. If $j_{\epsilon_{m-1}}(v^{-})\neq j$ then $b_{m}f=0$ and therefore $e=0$. Otherwise we have that $b_{m}f=e(v^{-})\otimes e(\epsilon_{m-1},i)$ with $i\neq j_{\epsilon_{m-1}}(v^{-})$. Applying ˜6.2 to $e=b_{m}f$ and $\alpha=\alpha_{m}$ we obtain the expected result. Now if $\alpha_{m}=0$, knowing that $\alpha_{m-1}\neq 0$ we find $v\in\mathcal{V}$ of length $p=\sum|\alpha_{i}|$ such that we have $u^{\alpha_{m-1}}b_{m-1}\cdots u^{\alpha_{0}}e(k)=\epsilon e(v)$. Applying $b_{m}$ to $e(v)$ we either find 0 or change the last coordinate $i_{p}$ of $v$.
We now prove the freeness with amalgamation of $u$ and $\mathcal{M}$, i.e. that (17) holds true for all $e\in\Xi$ and for all $b_{i}=\Pi(B_{i})$ where $B_{i}\in\mathrm{M}_{N}(\mathbb{C})$ and $\mathrm{E}(B_{i})=0$. Owing to the first assertion of ˜6.1, we only have to check (17) holds for $e\in H_{\emptyset}$ basis vector. Finally by linearity in $b_{i}$ for all $i\leq n$ it is sufficient to consider that $B_{i}$ is an elementary matrix $E_{k,\ell}$ of $\mathrm{M}_{N}(\mathbb{C})$ with zeros on the diagonal. Now let $k\in[N]$, $m\geq 2$, $\alpha_{0},...,\alpha_{m}\in\mathbb{Z}$ such that $\alpha_{i}\neq 0$ for all $i\notin\{0,m\}$ and $b_{1},...,b_{m}$ such that $b_{i}=\Pi(B_{i})$ where $B_{i}$ elementary matrix. Owing to the second assertion in ˜6.1, we have $f=u^{\alpha_{m}}b_{m}\cdots u^{\alpha_{1}}b_{1}u^{\alpha_{0}}e(k)=\epsilon e(v)$ for some $v\in\mathcal{V}$ such that $|v|\geq 1$. In particular $f$ is indeed orthogonal to $e$. ∎
All the ingredients are now gathered to prove ˜2.1.
Proof of ˜2.1.
From ˜6.1 and (16) from ˜6.2, for fixed non-all zero integers $\ell_{0},...,\ell_{p}$ we have
$\operatorname{Tr}_{N}(Q^{\ell_{0}}{Q^{\ast}}^{\ell_{1}}\cdots Q^{\ell_{p-1}}{Q^{\ast}}^{\ell_{p}})-\Phi({q}^{\ell_{0}}{{q}^{\ast}}^{\ell_{1}}\cdots{q}^{\ell_{p-1}}{{q}^{\ast}}^{\ell_{p}})=\frac{\delta^{\sum_{t}\ell_{t}}}{N}\underset{N\rightarrow\infty}{\rightarrow}0.$
Besides, as mentioned above the freeness with amalgamation is a direct consequence of (17). ∎
@@ 6.4 Spectral radius of the limit operator
$\mathcal{L}(T,n):=\left\{\vec{n}=(n_{1},...,n_{T})\in\{0,...,n\}^{T}:\penalty 10000\ \sum_{t=1}^{T}n_{t}+T-1=n\right\}.$
We first express $\Phi[\mathrm{b}_{\mathrm{x}}^{n_{1}}(\mathrm{b}_{\mathrm{x}}^{*})^{n_{2}}]$ for any integer $n_{1},n_{2}\geq 0$ and $\mathrm{b}_{\mathrm{x}}$ defined by (9).
@@
$\Phi[\mathrm{b}_{\mathrm{s}}^{n_{1}}(\mathrm{b}_{\mathrm{s}}^{*})^{n_{2}}]=\sum_{T=1}^{n_{1}\land n_{2}}\sum_{\begin{subarray}{c}\vec{n}\in\mathcal{L}(T,n_{1})\\ \penalty 10000\ \vec{n}^{\prime}\in\mathcal{L}(T,n_{2})\penalty 10000\ \end{subarray}}\prod_{t=1}^{T}\operatorname{Tr}_{N}\left(Q^{n_{t}}(Q^{*})^{n_{t}^{\prime}}\right)$
@@
$\Phi[\mathrm{b}_{\mathrm{c}}^{n_{1}}(\mathrm{b}_{\mathrm{c}}^{*})^{n_{2}}]=\sum_{T=1}^{n_{1}\land n_{2}}\sum_{\begin{subarray}{c}\vec{n}\in\mathcal{L}(T,n_{1})\\ \penalty 10000\ \vec{n}^{\prime}\in\mathcal{L}(T,n_{2})\penalty 10000\ \end{subarray}}(1-r)^{2(T-1)}\prod_{t=1}^{T}\operatorname{Tr}_{N}\left(\{rQ\}^{n_{t}}\{rQ^{*}\}^{n_{t}^{\prime}}\right).$
@@
$\mathcal{F}_{t}:=\sigma\left(\{\alpha\},\{j_{\epsilon}(s)\penalty 10000\ ;s\in\mathcal{S},\penalty 10000\ |s|\leq t-1,\penalty 10000\ \epsilon\in\{+,-\}\}\right).$
Then Section˜6.4 is a direct consequence of the following lemma.
@@
According to the first assertion of the lemma, we find a unique $(w_{T-1},w^{\prime}_{T-1})\in\mathcal{V}^{2}$ of length T1T-1 such that $e_{T-1}=e(w_{T-1})$ and $e^{\prime}_{T-1}=e(w_{T-1}^{\prime})$. Then rewriting $I_{T}$ and supposing that $w_{T-1}$ and $w_{T-1}^{\prime}$ are non-zero and therefore $\epsilon(w_{T-1})=\epsilon(w_{T-1}^{\prime})=+$ we obtain
@@
where for $M\in\mathrm{M}_{N}(\mathbb{C})$ we use the notation $(M)_{k}:=(M)_{k,k}$. Noticing that for all $w\in\mathcal{V}$ such that |w|=T1|w|=T-1 we have that $j_{+}(w)$ is independent from $\mathcal{F}_{T-1}$ we have
$\mathbb{E}(I_{T}|\mathcal{F}_{T-1})=\sum_{w\in\mathcal{V}\penalty 10000\ ;|w|=T-1}\mathbb{E}((b_{T}c_{T})_{j_{+}(w)})\mathds{1}(w_{T-1}=w_{T-1}^{\prime}=w)=\operatorname{Tr}_{N}(b_{T}c_{T})\langle e(w_{T-1}),e(w_{T-1}^{\prime})\rangle.$
@@
For all $B\in\mathrm{M}_{N}(\mathbb{C})$ and $n\geq 1$ integer we have
@@
Let $e\in\mathrm{H}$ such that $\|e\|_{2}=1$. We set $b:=\Pi(B)$. We have
$u+b]^{n}(e)=u^{n}(e)+\sum_{T=0}^{n}\sum_{\sum_{t=1}^{T}\lambda_{t}=n-T+1}b^{\lambda_{1}}ub^{\lambda_{2}}u\cdots ub^{\lambda_{T}}$
where in the term $b^{\lambda_{1}}ub^{\lambda_{2}}u\cdots ub^{\lambda_{T}}$, the number of occurrences of $u$ is bounded by n1n-1, therefore the sum is orthogonal to $u^{n}(e)$. It remains to take the norm of in the equality above. ∎
7 Path decomposition and computation on random permutation
7.1 Main technical statement
Having proved the first statement of ˜2.1 about the existence of the operator $(q,u)\in\mathrm{B}(\mathrm{H})$, the second statement is proved by applying the following theorem. First note that for any $\mathrm{x}\in\{\mathrm{s},\mathrm{c}\}$ and $\ell\geq 0$ integer we have
$\left|\lambda_{2}(A_{\mathrm{x}})\right|=\{|\lambda_{2}(A_{\mathrm{x}}^{\ell})|\}^{1/\ell}\leq\|A_{\mathrm{x}}|_{\boldsymbol{1}^{\perp}}^{\ell}\|^{1/\ell}$ (19)
where we used the fact that $\boldsymbol{1}$ is eigenvector for $A_{\mathrm{x}}$ and $A_{\mathrm{x}}^{\ast}$ and therefore $\lambda_{2}(A_{\mathrm{x}})=\lambda_{1}(A_{\mathrm{x}}|_{\boldsymbol{1}^{\perp}})$. To simplify further computations, for any integer $\ell\geq 0$ we introduce
$A_{\ell,\mathrm{x}}:=MA_{\mathrm{x}}^{\ell}.$
Since $M$ is a permutation and $Q$ verifies (H:1), $A_{\ell,\mathrm{x}}$ leaves $\boldsymbol{1}$ and $\boldsymbol{1}^{\perp}$ stable and the multiplication by $M$ above does not change the operator norm. Furthermore for $N$ large enough we have
@@ 7.1 Main technical statement
Where $\epsilon=\epsilon(N)$ is given by ˜2.1. Therefore for $N$ large enough
$\mathbb{P}\left(\|A_{\mathrm{x}}^{\ell}|_{\boldsymbol{1}^{\perp}}\|\geq(1+\epsilon)^{\ell}\rho_{n,\mathrm{x}}^{\ell}\right)=\mathbb{P}\left(\|A_{\ell,\mathrm{x}}|_{\mathbf{1}^{\perp}}\|\geq(1+\epsilon)^{\ell}\rho_{n,\mathrm{x}}^{\ell}\right)\leq\mathbb{P}\left(\|{A_{\ell,\mathrm{x}}|}_{\boldsymbol{1}^{\perp}}\|\geq e^{\frac{\epsilon\ell}{2}}\rho_{n,\mathrm{x}}^{\ell}\right).$
Adjusting the constants, the second statement of ˜2.1 becomes a corollary of the following result applied to $\ell=\frac{(1-c)\ln(N)}{7K_{\mathrm{x}}\ln(d)}$ where
$K_{\mathrm{s}}=1\vee(2\ln(1+\delta+\kappa))\penalty 10000\ \penalty 10000\ \text{and}\penalty 10000\ \penalty 10000\ K_{\mathrm{c}}=1\vee\left(2\ln\left(\frac{1+r\delta+r\kappa}{1-r}\right)\right).$ (20)
The condition (H:2) implies that $K_{\mathrm{x}}^{\prime}$ is uniformly bounded in $N$.
@@
Let $n\geq 0$, $(M,Q)$ and $\rho_{n,\mathrm{x}}$ be as in ˜2.1. For any $0<c<1$, there exists $C>0$ depending on $n$, $K$ and $\alpha$, such that for any $1\leq\ell\leq\frac{\ln(N)}{7\ln(d)}$ we have
$\mathbb{P}\left(\|A_{\mathrm{x}}^{\ell}|_{\boldsymbol{1}^{\perp}}\|\geq e^{C\ln(N)^{1-\alpha}}\rho_{n,\mathrm{x}}^{\ell}\right)\leq N^{-c}+C\frac{\ell^{4}(4d)^{3\ell}}{N}.$
To estimate an upper bound on $\|(A_{\ell,\mathrm{x}})|_{\boldsymbol{1}^{\perp}}\|$, we use the high trace method, i.e. Markov inequality combined with the inequality